12. Convert to polar coordinates and evaluate: 1 LL62 +7) dydx + y2 +
Answer #1

Given intergal dyda 2-1 (2+y+1) y=11-28 Limits се lo Y = -r-got to yo ri- Let 2= cose y = visino Since y=ri-20 osin o re-valo(ari) (In(a) & IT In(a) O. 6930 2:177 28 not yety dyda e. Ti lua) -1 -Tige